# Spectraformer: A Unified Random Feature Framework for Transformer

### v3

@@
Abstract
Linearization of attention using various kernel approximation and kernel learning techniques has shown promise. Past methods use a subset of combinations of component functions and weight matrices within the random features paradigm. We identify the need for a systematic comparison of different combinations of weight matrices and component functions for attention learning in Transformer. In this work, we introduce Spectraformer, a unified framework for approximating and learning the kernel function in linearized attention of the Transformer. We experiment with broad classes of component functions and weight matrices for three textual tasks in the LRA benchmark. Our empirical findings indicate that different kernels are good at different tasks and that kernel choice is fundamental to performant models. Our code is available at: https://anonymous.4open.science/r/spectraformer-8A97.
1 Introduction
@@
Transformer (Vaswani et al., 2017) has revolutionized the landscape of natural language processing (NLP) and forms the basis of almost all state-of-the-art language models. Its influence has reached beyond NLP into computer vision (Han et al., 2023), speech processing (Lin et al., 2022) and other fields. Compared to its precursors like the LSTMs, Transformer leverages parallelism since it is fully based on the attention mechanism. Therefore, the performance of Transformer is tied to the effective use of attention. Attention in the Original Transformer (referred to as ‘OT’ hereafter) uses the softmax function which is quadratic in time complexity. However, softmax is not fundamental to attention. The softmax serves the role of a compatibility function, which captures the degree of compatibility between two tokens (see Equations 1 and 2). Extensive experiments with other compatibility functions show that the softmax is not the only possibility, and that the ‘best’ compatibility function depends on the task (Tsai et al., 2019).
@@
Replacing the softmax formulation with alternatives may allow us to reduce the time complexity of attention computation. In this paper, we focus on the subset of alternatives that employs kernel functions. Kernel functions have been historically used in machine learning algorithms to simplify the estimation of parameters (Hofmann et al., 2008). Tsai et al. (2019) show that attention can be formulated using kernels and then be linearized (Katharopoulos et al., 2020), given the feature map of the respective attention kernel. Random feature-based algorithms (Liu et al., 2022) are a family of algorithms, inspired by spectral analysis, which provides the feature map associated with a kernel. A random feature consists of three components: component function, weight matrix, and component weight (see Section 3). Due to their robustness and flexibility, random features give rise to ‘random feature Transformers’, the family of Transformers with linearized attention via random features (hereby abbreviated as RF), first with the work by Choromanski et al. (2021). This is seen to lead to three strands of research: (A) Optimization of weight matrices either by making the matrices have lower time or space complexity, or by incorporating beneficial properties like orthogonality to reduce variance (Yu et al., 2016; Choromanski et al., 2021; 2022; Reid et al., 2023); (B) Enhancement of component functions by either: (i) finding a function with a tighter variance or approximation bound, or (ii) be engineered to output with numerical stability and be bounded (Choromanski et al., 2021; Likhosherstov et al., 2022; 2023); (C) Parameterization of weight matrices to perform kernel learning instead of kernel approximation (Chowdhury et al., 2022). (A) and (B) seek to improve approximation quality via approximation error and variance reduction. (C) dispenses with the approximation of the attention kernel (typically the softmax), and makes the kernel itself learnable, to attain a better performance. However, the three have been explored separately, creating a situation where there are different ways to construct the random features. Additionally, past works only explore certain combinations of component functions and weight matrices, creating disjoint overlaps and gaps in the literature. Finally, the weight matrices are often those explicitly designed in the context of kernelized attention in the Transformer (Yu et al., 2016; Choromanski et al., 2021; 2022; Reid et al., 2023), ignoring those which are popular in the kernel methods literature (e.g., Liu et al. (2022)). These limitations necessitate a unified framework that allows the comparison of different combinations in search of the best-performing one.
We introduce Spectraformer, a unified framework of random feature-based attention in Transformer. It utilizes spectral analysis-inspired random features to construct linearized attention in Transformers. Spectraformer allows for the experimentation of various combinations of weight matrices and component functions. We will open-source our codebase upon acceptance. Through our benchmarking experiments on three textual Long Range Arena (Tay et al., 2021) tasks and 18 combinations of weight matrices and component functions, Spectraformer enables novel combinations of component functions and weight matrices from the kernel literature to be experimented easily in the Transformer setting. This process makes possible the discovery of combinations that outperform existing random feature-based Transformers in terms of performance, training time, and peak memory consumption. Various novel Spectraformer combinations are more accurate than the previous SOTA models, SADERF-ORF (a.k.a., FAVOR# (Likhosherstov et al., 2023)), OPRF-ORF (a.k.a., FAVOR++ (Likhosherstov et al., 2022)), PosRF-ORF (a.k.a., FAVOR+ (Choromanski et al., 2021)), whilst significantly reducing training time and memory consumption over both. We hope that Spectraformer will enable the mechanism to benchmark more random feature-based Transformers for these and other tasks. Our contributions are:
  1. 1.
@@
    Our novel framework, Spectraformer, which generalizes over all past works in linearized attention with kernel approximation and kernel learning.
  2. 2.
    We demonstrate how Spectraformer can discover novel SOTA models in the random feature Transformer family for the LRA benchmark and that these models are more performant than existing random feature Transformers.
  3. 3.
    We empirically identify the relationship between performant combinations and task characteristics, aiding the right kernel choice when tackling a task.
  4. 4.
    Our code is currently available for review, but they will be publicly released upon acceptance. We detail guide on how to incorporate future kernel works into Spectraformer. With this work, every future kernel work that is introduced can easily be incorporated into the transformer.
2 Preliminaries
We first cover the theoretical foundations to this work. A list of commonly used acronyms is available in Table 1.
Table 1: Table of acronyms
Acronym Description
OT (Vaswani et al., 2017) Original Transformer
RF (Rahimi & Recht, 2007) Random Feature method
RBF Radial Basis Function kernel (also called the Gaussian kernel)
TrigRF (Rahimi & Recht, 2007) Trigonometric RF (component function)
PosRF (Choromanski et al., 2021) Positive RF (component function)
OPRF (Likhosherstov et al., 2022) Optimized positive RF (component function)
SADERF (Likhosherstov et al., 2023) Simplified Asymmetric Dense-Exponential RF (component function)
ORF (Yu et al., 2016) Orthogonal RF (weight matrix)
SORF (Yu et al., 2016) Structural ORF (weight matrix)
QMC (Avron et al., 2016) Quasi- Monte Carlo (weight matrix)
MM (Shen et al., 2017) (Liu et al., 2022) Moment Matching (weight matrix)
SGQ (Dao et al., 2017) Sparse Grid Quadrature (weight matrix)
FastFoodL(Yang et al., 2015) FastFood Learnable (weight matrix)
2.1 Kernelized Attention
@@
$\begin{array}[]{cccc}\bm{A}_{i}=\sum_{j=1}^{N}\frac{f(\bm{q}_{i},\bm{k}_{j})}{\sum_{l}f(\bm{q}_{i},\bm{k}_{l})}\bm{v}_{j}=\sum_{j=1}^{N}a_{j}\bm{v}_{j};&a=g(e);&e=f(\bm{q}_{i},\bm{k}_{j})\end{array}$ (1)
@@
$\begin{array}[]{cc}f(\bm{q}_{i},\textbf{K})=\frac{\bm{q}_{i}^{T}\textbf{K}}{\sqrt{d_{k}}}&g(\bm{x}_{j})=\frac{\exp(\bm{x}_{j})}{\sum_{l}\exp(\bm{x}_{l})}\end{array}$ (2)
@@
$\begin{array}[]{cc}\bm{A}_{i}=\sum_{j=1}^{N}\frac{\exp(\bm{q}_{i}\bm{k}_{j}^{T}/\sqrt{d_{k}})}{\sum_{l}\exp(\bm{q}_{i}\bm{k}_{l}^{T}/\sqrt{d_{k}})}\bm{v}_{j}&A=softmax(\frac{\bm{QK}^{T}}{\sqrt{d_{k}}})\bm{V}\end{array}$ (3)
@@
$\bm{A}_{i}=\sum_{j=1}^{N}\frac{\mathcal{K}(\bm{q}_{i},\bm{k}_{j})}{\sum_{l}\mathcal{K}(\bm{q}_{i},\bm{k}_{l})}\bm{v}_{j}$ (4)
As shown in Tsai et al. (2019), this is indeed the Nadaraya-Watson Kernel Estimator, with $\mathbb{E}_{p(\bm{k}_{j}|\bm{q}_{i})}[\bm{v}_{j}|X=\bm{k}_{j}]=\bm{A}_{i}$, $l_{i}\left(\bm{k}_{j})=p(\bm{k}_{j}|\bm{q}_{i}\right)$, $Y_{i}=\bm{v}_{j}$:
$\begin{array}[]{ccc}\mathbb{E}_{l_{i}}[Y_{i}|X=\bm{u}]=\sum_{i=1}^{N}l_{i}(\bm{u})Y_{i};&l_{i}(\bm{u})=\frac{\mathcal{K}(\frac{\bm{u}-\bm{x}_{i}}{h})}{\sum_{j=1}^{n}\mathcal{K}(\frac{\bm{u}-\bm{x}_{j}}{h})}\end{array}$ (5)
@@
$\mathcal{K}_{softmax}(\bm{q}_{i},\bm{k}_{j})=\exp(\bm{q}_{i}^{T}\bm{k}_{j}/\sqrt{d_{k}})$ (6)
2.2 Linearizing Kernelized Attention via Random Features

### v5

@@
Abstract.
Linearization of attention using various kernel approximation and kernel learning techniques has shown promise. Past methods used a subset of combinations of component functions and weight matrices within the random feature paradigm. We identify the need for a systematic comparison of different combinations of weight matrices and component functions for attention learning in Transformer. Hence, we introduce Spectraformer, a unified framework for approximating and learning the kernel function in the attention mechanism of the Transformer. Our empirical results demonstrate, for the first time, that a random feature-based approach can achieve performance comparable to top-performing sparse and low-rank methods on the challenging Long Range Arena benchmark. Thus, we establish a new state-of-the-art for random feature-based efficient Transformers. The framework also produces many variants that offer different advantages in accuracy, training time, and memory consumption. Our code is available at: https://github.com/cruiseresearchgroup/spectraformer.
transformers, kernel, linearized attention, kernelized attention
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JACMccs: Computing methodologies Kernel methodsccs: Computing methodologies Neural networksccs: Computing methodologies Natural language processing
1. Introduction
@@
Transformer (Vaswani et al., 2017) has revolutionized the landscape of natural language processing (NLP) and forms the basis of almost all state-of-the-art language models. Its influence has reached beyond NLP into computer vision (Han et al., 2023; Yu et al., 2024; Yuan et al., 2025; Hao et al., 2025), speech processing (Lin et al., 2022), and other fields. Compared to its precursors like the LSTMs, Transformer leverages parallelism since it is fully based on the attention mechanism. Therefore, the performance of Transformer is tied to the effective use of attention. Attention in the Original Transformer (referred to as ‘OT’ hereafter) uses the softmax function which is quadratic in time complexity.
To mitigate the quadratic complexity of the standard attention mechanism, a diverse range of efficient Transformer architectures have been developed (Tay et al., 2022). These include memory/ downsampling methods (Big Bird (Zaheer et al., 2020), Nystromformer (Xiong et al., 2021), and Skyformer (Chen et al., 2021)), learnable pattern methods (Reformer (Kitaev et al., 2020)), Informer (Zhou et al., 2021), Linformer (Wang et al., 2020). Our work, Spectraformer, focuses on another significant category: the linearization of attention through a unified framework of random features. We approach this linearization by rethinking the role of the softmax. The softmax acts as a compatibility function, capturing the similarity between pairs of tokens (see Equations 1 and 2). Extensive experiments with other compatibility functions show that the softmax is not the only possibility, and that the ‘best’ compatibility function depends on the task (Tsai et al., 2019).
@@
Replacing the softmax formulation with alternatives may allow us to reduce the time complexity of attention computation. In this paper, we focus on the subset of alternatives that employs kernel functions. Kernel functions have been historically used in machine learning algorithms to simplify the estimation of parameters (Hofmann et al., 2008). Tsai et al. (2019) show that attention can be formulated using kernels and then linearized (Katharopoulos et al., 2020), given the feature map of the respective attention kernel. Random feature-based algorithms (Liu et al., 2022) are a family of algorithms, inspired by spectral analysis, which provides the feature map associated with a kernel. A random feature consists of three components: component function, weight matrix, and component weight (see Section 3). Due to their robustness and flexibility, random features give rise to ‘random feature Transformers’, the family of Transformers with linearized attention via random features (hereby abbreviated as RF), first with the work by Choromanski et al. (2021). This is seen to lead to three strands of research: (A) Optimization of weight matrices either (i) by making the matrices have lower time or space complexity, or (ii) by incorporating beneficial properties like orthogonality to reduce variance (Yu et al., 2016; Choromanski et al., 2021, 2022; Reid et al., 2023); (B) Enhancement of component functions by either: (i) finding a function with a tighter variance or approximation bound, or (ii) engineering in order to output with numerical stability and be bounded (Choromanski et al., 2021; Likhosherstov et al., 2022, 2023); (C) Parameterization of weight matrices to perform kernel learning instead of kernel approximation (Chowdhury et al., 2022). (A) and (B) seek to improve approximation quality via approximation error and variance reduction. (C) dispenses with the approximation of the attention kernel (typically the softmax), and makes the kernel itself learnable, to attain a better performance. However, the three have been explored separately, creating a situation where there are different ways to construct the random features. Additionally, past works only explore certain combinations of component functions and weight matrices, creating disjoint overlaps and gaps in the literature. Finally, the weight matrices are often those explicitly designed in the context of kernelized attention in the Transformer (Yu et al., 2016; Choromanski et al., 2021, 2022; Reid et al., 2023), ignoring those which are popular in the kernel method literature (e.g., Liu et al. (2022)). These limitations highlight the need for a unified framework to systematically compare combinations and identify optimal configurations.
Table 1. Experimental results of Spectraformer variants against previous SOTA efficient Transformers on the LRA benchmark. All models were run for five seeds using 128 random features. We report mean accuracy (on the test set), mean time (training time (hour)), and mean memory (peak memory consumption (GB)). $L$: ListOps, $T$: Text, $R$: Retrieval, II: Image, $P$: Pathfinder, and $\mu$ for the average across tasks. Entries are sorted in descending order by $\mu$. Best models per category are boldened.
Accuracy (%) $\uparrow$ Time (hour) $\downarrow$ Memory (GB) $\downarrow$
L T R I P $\mu$ L T R I P $\mu$ L T R I P $\mu$
Nystromformer 38.69 (0.59) 61.57 (0.38) 80.57 (0.30) 39.49 (0.89) 69.96 (1.27) 58.06 0.56 0.90 1.00 2.19 1.19 1.17 1.10 1.87 1.84 5.55 2.79 2.63
Big Bird 38.82 (0.52) 61.04 (0.15) 80.53 (0.32) 37.59 (1.04) 68.28 (6.74) 57.25 1.60 3.17 3.24 5.72 2.89 3.32 2.24 4.57 3.80 8.42 4.22 4.65
OPRF-FastFoodL 37.73 (0.28) 64.32 (0.61) 78.65 (1.94) 38.41 (0.66) 66.76 (0.39) 57.17 1.07 2.07 2.11 4.02 2.05 2.26 0.84 1.68 1.64 3.26 1.68 1.82
Reformer 35.50 (4.09) 61.28 (0.45) 78.31 (0.29) 43.62 (0.81) 66.28 (2.35) 57.00 0.65 1.29 1.29 2.51 1.26 1.40 1.61 3.20 2.95 6.38 3.20 3.47
Skyformer 38.96 (0.63) 60.67 (0.45) 81.90 (0.37) 32.93 (0.39) 69.81 (1.38) 56.85 0.78 1.35 1.49 3.09 1.63 1.67 1.67 3.01 2.92 7.90 3.96 3.89
OT 38.49 (0.80) 59.72 (1.07) 80.86 (0.20) 37.48 (0.83) 67.56 (6.12) 56.82 1.96 7.27 7.07 4.30 2.20 4.56 4.37 16.72 8.74 9.42 4.72 8.79
OPRF-SGQ 37.08 (0.56) 61.09 (0.77) 79.39 (0.94) 36.68 (0.28) 67.53 (2.79) 56.35 0.68 1.27 1.25 2.46 1.29 1.39 1.33 2.64 2.50 5.25 2.64 2.87
PosRF-$M$ 37.13 (0.27) 62.88 (1.17) 80.58 (0.44) 33.82 (0.51) 67.11 (0.45) 56.30 0.56 1.05 1.06 2.02 1.05 1.15 1.14 2.26 2.05 4.49 2.26 2.44
OPRF-OR 38.05 (0.78) 60.18 (0.49) 81.19 (0.23) 33.70 (0.57) 67.24 (0.88) 56.07 0.68 1.26 1.24 2.47 1.29 1.39 1.33 2.64 2.50 5.25 2.64 2.87
SADERF-QMC 37.18 (0.47) 60.61 (2.06) 80.67 (0.44) 34.55 (0.63) 67.32 (0.40) 56.07 0.68 1.24 1.28 2.47 1.26 1.39 1.40 2.79 2.63 5.56 2.79 3.04
PosRF-QMC 37.05 (0.16) 60.93 (0.59) 80.67 (0.29) 34.03 (0.94) 67.25 (0.44) 55.99 0.55 1.05 1.05 1.99 1.05 1.14 1.14 2.26 2.05 4.49 2.26 2.44
OPRF-QMC 37.86 (0.40) 60.87 (1.85) 80.44 (0.20) 33.87 (0.69) 66.81 (0.68) 55.97 0.68 1.26 1.24 2.43 1.28 1.38 1.33 2.64 2.50 5.25 2.64 2.87
SADERF-ORF 37.41 (0.45) 59.92 (0.90) 81.05 (0.18) 33.06 (1.02) 67.11 (0.56) 55.71 0.69 1.25 1.28 2.50 1.28 1.40 1.40 2.79 2.63 5.56 2.79 3.04
OPRF-$M$ 38.31 (0.38) 60.01 (1.01) 81.03 (0.31) 33.93 (0.85) 65.17 (5.17) 55.69 0.68 1.26 1.29 2.47 1.28 1.40 1.33 2.64 2.50 5.25 2.64 2.87
SADERF-$M$ 37.17 (0.27) 60.81 (1.82) 81.05 (0.20) 34.03 (1.04) 65.31 (4.85) 55.67 0.69 1.25 1.28 2.49 1.28 1.40 1.40 2.79 2.63 5.56 2.79 3.04
SADERF-SGQ 37.22 (0.28) 62.86 (0.96) 78.51 (0.68) 37.82 (0.59) 61.27 (5.38) 55.54 0.68 1.25 1.28 2.50 1.27 1.39 1.40 2.79 2.63 5.56 2.79 3.04
SADERF-FastFoodL 29.72 (7.27) 64.71 (0.45) 77.50 (0.96) 38.38 (0.82) 66.47 (1.47) 55.35 1.09 2.09 2.18 4.05 2.06 2.29 0.90 1.80 1.76 3.52 1.80 1.96
PosRF-FastFoodL 29.54 (5.46) 64.61 (0.29) 77.10 (1.02) 38.28 (0.43) 66.38 (0.71) 55.18 1.02 2.00 2.01 3.88 1.98 2.18 0.77 1.54 1.50 3.01 1.54 1.67
PosRF-ORF 34.50 (6.49) 61.10 (1.76) 80.53 (0.33) 33.72 (1.03) 65.52 (4.89) 55.07 0.56 1.05 1.06 2.02 1.06 1.15 1.14 2.26 2.05 4.49 2.26 2.44
OPRF-SORF 29.59 (3.77) 64.81 (0.22) 77.12 (0.78) 37.42 (0.47) 63.98 (5.38) 54.58 0.68 1.26 1.24 2.42 1.28 1.38 1.33 2.64 2.50 5.25 2.64 2.87
SADERF-SORF 34.31 (2.50) 64.82 (0.26) 75.24 (1.12) 36.76 (1.22) 60.98 (4.90) 54.42 0.68 1.25 1.28 2.47 1.26 1.39 1.40 2.79 2.63 5.56 2.79 3.04
Linformer 36.94 (0.60) 56.99 (1.23) 77.86 (0.35) 39.09 (0.78) 59.24 (5.84) 54.02 0.41 0.74 0.77 1.28 0.71 0.78 0.86 1.70 1.61 3.39 1.71 1.85
PosRF-SGQ 28.81 (7.77) 62.30 (0.47) 78.31 (0.39) 37.96 (0.76) 59.71 (7.87) 53.42 0.56 1.06 1.05 2.01 1.06 1.15 1.14 2.26 2.05 4.49 2.26 2.44
Informer 31.45 (7.20) 62.06 (1.20) 76.92 (0.14) 37.90 (0.30) 57.07 (6.57) 53.08 0.88 1.92 2.23 3.21 1.78 2.01 2.66 5.57 3.27 5.20 2.61 3.86
PosRF-SORF 21.91 (5.98) 62.06 (1.36) 66.33 (0.41) 28.70 (3.87) 52.37 (5.24) 46.27 0.55 1.06 1.06 2.00 1.05 1.14 1.14 2.26 2.05 4.49 2.26 2.44
We introduce Spectraformer, a unified framework for approximating and learning the kernel function in the attention mechanism of the Transformer through the lens of random features. While other methods (e.g., Nystromformer (Xiong et al., 2021), BigBird (Zaheer et al., 2020)) have explored alternative approaches to improving efficiency, Spectraformer provides a systematic way to explore the vast design space of random feature-based attention. It utilizes spectral analysis-inspired random features to construct linearized attention in Transformers. Spectraformer allows for the experimentation of various combinations of weight matrices and component functions. Through our benchmarking experiments on the Long Range Arena (LRA) (Tay et al., 2021) and 18 combinations of weight matrices and component functions, Spectraformer produces novel combinations of component functions and weight matrices that perform competitively against previous efficient Transformers with different computational advantages. In addition, our work demonstrates that a well-configured kernelized attention model can perform on the same level as other classes of efficient Transformers, with our best Spectraformer variant proving highly competitive with leading methods like Nystromformer (Xiong et al., 2021) and BigBird (Zaheer et al., 2020). In doing so, it establishes a new state-of-the-art for random feature-based methods with a mean accuracy improvement of 0.35% over the Original Transformer (OT). Many Spectraformer variants also outperform existing efficient Transformers with random feature attention such as SADERF-ORF (a.k.a., FAVOR# (Likhosherstov et al., 2023)), OPRF-ORF (a.k.a., FAVOR++ (Likhosherstov et al., 2022)), PosRF-ORF (a.k.a., FAVOR+ (Choromanski et al., 2021)). Our contributions are:
  1. (1)
@@
    A novel framework, Spectraformer, that unifies and generalizes past work on linearized attention using both kernel approximation and kernel learning.
  2. (2)
    Establishment of a new state-of-the-art for random feature-based efficient Transformers, demonstrating that this model class can be highly competitive with leading sparse and low-rank methods on the Long Range Arena benchmark.
  3. (3)
    A comprehensive empirical analysis of 18 combinations of component functions and weight matrices, offering practical guidance on selecting configurations based on trade-offs between accuracy, memory, and training time.
  4. (4)
    A public and extensible implementation designed to facilitate future research by allowing for the easy integration of new kernel-based techniques into Transformers.
2. Preliminaries
We first cover the theoretical foundations of this work.
2.1. Kernelized Attention
@@
(1) $\begin{array}[]{cccc}\bm{A}_{i}=\sum_{j=1}^{N}\frac{f(\bm{q}_{i},\bm{k}_{j})}{\sum_{l}f(\bm{q}_{i},\bm{k}_{l})}\bm{v}_{j}=\sum_{j=1}^{N}a_{j}\bm{v}_{j};&a=g(e);&e=f(\bm{q}_{i},\bm{k}_{j})\end{array}$
@@
(2) $\begin{array}[]{cc}f(\bm{q}_{i},\textbf{K})=\frac{\bm{q}_{i}^{T}\textbf{K}}{\sqrt{d_{k}}}&g(\bm{x}_{j})=\frac{\exp(\bm{x}_{j})}{\sum_{l}\exp(\bm{x}_{l})}\end{array}$
@@
(3) $\begin{array}[]{cc}\bm{A}_{i}=\sum_{j=1}^{N}\frac{\exp(\bm{q}_{i}\bm{k}_{j}^{T}/\sqrt{d_{k}})}{\sum_{l}\exp(\bm{q}_{i}\bm{k}_{l}^{T}/\sqrt{d_{k}})}\bm{v}_{j}&A=softmax(\frac{\bm{QK}^{T}}{\sqrt{d_{k}}})\bm{V}\end{array}$
@@
(4) $\bm{A}_{i}=\sum_{j=1}^{N}\frac{\mathcal{K}(\bm{q}_{i},\bm{k}_{j})}{\sum_{l}\mathcal{K}(\bm{q}_{i},\bm{k}_{l})}\bm{v}_{j}$
As shown by Tsai et al. (2019), this is indeed the Nadaraya-Watson Kernel Estimator, with
$\mathbb{E}_{p(\bm{k}_{j}|\bm{q}_{i})}[\bm{v}_{j}|X=\bm{k}_{j}]=\bm{A}_{i}$, $l_{i}\left(\bm{k}_{j})=p(\bm{k}_{j}|\bm{q}_{i}\right)$, $Y_{i}=\bm{v}_{j}$:
(5) $\begin{array}[]{ccc}\mathbb{E}_{l_{i}}[Y_{i}|X=\bm{u}]=\sum_{i=1}^{N}l_{i}(\bm{u})Y_{i};&l_{i}(\bm{u})=\frac{\mathcal{K}(\frac{\bm{u}-\bm{x}_{i}}{h})}{\sum_{j=1}^{n}\mathcal{K}(\frac{\bm{u}-\bm{x}_{j}}{h})}\end{array}$
@@
(6) $\mathcal{K}_{softmax}(\bm{q}_{i},\bm{k}_{j})=\exp(\bm{q}_{i}^{T}\bm{k}_{j}/\sqrt{d_{k}})$
2.2. Linearizing Kernelized Attention via Random Features